For the following set of analyses of an urban air sample for carbon monoxide: 325, 320, 334, 331, 280, 331, 338 mg/m3

(b) calculate the coefficient of variation and RSD.

(c) is there any value in the above data set that can be discarded as an outlier?
Answer #1

mean = sum of terms/ total number of terms

= 325+320+334+331+280+331+338 /7

=2259/7

=322.71

S.D= 2.6901

coefficient of variation= S.D/mean

= 2.6901/322.71

=0.0083

RSD= COV×100%

= 0.0083×100%

=0.83%

C)- the value 280 is can be discarded as an outlier